# Threat Modeling for AI: The Case for an Asset-Centric Approach

### v1

@@ II-B Training
II-C Deploy & Inference
This stage describes the process for deploying a model, including serving inference requests, monitoring, and model maintenance. Common operations are outlined in Figure 5.
Types of assets, and their relationships, that exist within this stage include:
  • Model (Consumed and Exposed). Obtained from the training stage. The model is typically considered confidential, but its confidentiality can be impacted as adversaries can infer properties about it through their interactions. They can correlate the model behavior on inputs and outputs to steal or attack the model.
  • Inference or Prompt Inputs (Consumed, Modified) and Outputs (Produced). Inputs are processed with the model to generate corresponding outputs. The source of inputs is product specific. They might originate from users, sensors, or system calls. Input pre-processing might augment them with additional information; for example, from API calls or system state. Similarly, outputs may be transmitted to users, other AIs or services, etc. Conventional security best practices call for inputs and outputs to be secured in transit and during any pre-processing. It is increasingly common practice to collect inputs for potential use during AI retraining. Users typically expect that their inputs and outputs are isolated from those of other users; to maintain their privacy.
  • RAG Dataset (Consumed). Represents a dataset deployed alongside a large language model, used to augment user inputs with data deemed to be relevant. This dataset can contain confidential or sensitive data. Attackers could also influence or change outputs by affecting this dataset.
  • Performance Metrics (Produced). These metrics describe the model’s performance and behavior on real world data, potentially motivating changes in this and other stages. These metrics are often used to detect drift or decide which input data to collect and use for retraining. They can include user feedback provided on model responses.
  • Raw Data (Produced). Inputs which are collected for retraining are consumed as Raw Data by subsequent Dataset Collection & Assembly stages.
  • Training Dataset (Exposed). The training dataset, unlike a RAG dataset, is not deployed alongside the model. This prevents adversaries from directly stealing or compromising it. However, the training dataset is, to some extent, encoded into the AI because of memorization [13]. Adversaries can, therefore, characterize the model to infer properties about the dataset or to extract information about individual data samples [13, 11].
@@ II-C Deploy & Inference
II-D Noteworthy Security Implications: Influencing AI Outputs
There are various noteworthy ways in which adversaries can exploit the relationships introduced above to influence AI Outputs by compromising the dataset or model. An adversary could trick the model into learning false features by changing the dataset [10, 38] or by directly compromising the model itself [8, 23].
Compromising the Dataset.
  • Permanent Writes to the training data, at any point during Dataset Collection & Assembly or Training [19, 10, 34].
  • Permanent Additions to the training data, by injecting malicious data into the data collection pipelines within Dataset Collection & Assembly or Deploy & Inference [30].
  • Transient Writes, by modifying data as the model consumes it, during Training. For example, by compromising a library for Dataset Augmentation.
Compromising the Model.
  • Permanent Writes to the model, during Training or Deploy & Inference [8].
  • Transient Writes, such as fault injections [28, 17, 23, 12], during Deploy & Inference.
The subtle relationships between AI assets, stemming from AI’s computing paradigm, result in unique challenges for security analysis.
III Challenges to Overcome
@@
Challenge 2. Understanding the impact of a vulnerability requires a holistic evaluation. Distributing an AI’s implementation and dependencies distributes its assets accordingly. Consider all the processes outlined in Figures 34, and 5. Every process can exist within a separate hardware and software context, and each performs unique operations or transformations on a subset of the system’s AI Assets. This creates many, often disjoint, attack surfaces. Reasoning about all possible implications of a vulnerability – identified in isolation, within a single context – requires that defenders map out how that vulnerability can be chained with vulnerabilities across other contexts.
@@
Our methodology addresses these challenges using an asset-centric approach. Section IV provides background on asset-centric threat modeling, and Section V describes how we build on it.
IV Asset-Centric Threat Modeling
The goal of Asset-Centric threat modeling is to identify vulnerabilities which, if exploited, allow an adversary to compromise an Asset.
@@ IV Asset-Centric Threat Modeling
IV-A Asset-Centric Reasoning about Threats
Although some assets are inherently valuable, others are valuable because they protect, or serve as a stepping stone towards, other assets. Adversaries must typically compromise this second class of assets through individual exploits to move towards their true goal. It is useful, therefore, to identify the set of assets an adversary must compromise to achieve their goal. Including the minimum capabilities they must obtain over each asset.
@@ IV-A Asset-Centric Reasoning about Threats
  • Cryptographic Key (Asset): Read access (Capability, compromising confidentiality).
  • Encrypted PII Data (Asset): Read access (Capability, compromising confidentiality).
In practice and in threat models, the vulnerabilities which provide each capability can be reasoned about – and exploited – as individual steps in a chain of attacks. Section V will describe how we leverage this property to develop our methodology.
V Asset-Centric Threat Modeling for AI
@@
In phase one (3, line 3) defenders map $AF$ by identifying capabilities adversaries could obtain over the AI assets relevant to their $system$. Defenders reason about these capabilities based on whether they originate within product scope (Section V-A) or are inherited from consumed assets (Section V-B). Our analysis implements this reasoning as two steps, however, there is no strict dependence between them. They can be performed in parallel with each other and alongside existing security analysis.
Phase two performs threat analysis (2, line 2) to generate a knowledge base of attacks $KB$, and then identifies which threats from $KB$ are relevant based on $AF$ (4, line 4). Section V-C describes how threats are broken down into their asset-centric requirements and added into $KB$. Section V-D describes how those threats are contextualized within a product, using $AF$, to determine if, and how, they are feasible.
@@
Phase 1, where defenders perform security analysis to generate $AF$ for their $system$, is outlined in Algorithm 2 and described in Sections V-A and V-B.
V-A Identifying Adversarial Influence within Product Scope (Phase 1)
This step guides defenders through AI-centric analysis of vulnerabilities within product scope. During this process, defenders will reason about vulnerabilities from their existing security analysis but may also identify additional related vulnerabilities, such as those described in [24, 18]. As outlined in 4, line 4-line 8, this step guides defenders through identifying the AI assets that exist within each in-scope stage, identifying vulnerabilities which could impact each asset, quantifying the capabilities each vulnerability could provide, and aggregating them into $AF$.
Repeataing this process within each stage – i.e. each system boundary – allows $AF$ to describe capabilities which could be obtained across across the entire development and deployment pipelines. This allows defenders to address Challenge 1 (Section III). An example output from this step is depicted on the top right of Figure 6.
V-A1 Holistic Evaluation
@@ V-A2 Enabling Collaboration
V-A3 Reasoning about Different Adversaries
@@
V-B Quantifying Assumptions about Adversarial Influence Outside Product Scope (Phase 1)
@@
V-C Threat Analysis (Phase 2)
The first step in phase two (2, line 2) generates $KB$ – a prerequisite for mapping attacks (4, line 4). Defenders repeat the first portion of the procedure in 15, line 15 for every attack they wish to add into $KB$. Fortunately, this process is product-independent. $KB$ can be reused when analyzing other products because it is not dependent on $system$. The cards on the bottom left of Figure 6 represent a knowledge base of threats.
As described in Section IV-A and depicted in 16, line 16, an attack’s requirements can be described using the set of assets an adversary must gain capabilities over. Papernot et. al.’s work in [26] presented a first step towards describing adversarial AI attacks based on their requirements; albeit less rigorously than would be necessary for bottom-up analysis. Section V-C1 describes the process for mapping the requirements for Backdoor attacks. Table I summarizes the identified requirements.
V-C1 Example Breakdown: Backdoor attacks
@@
    Manipulate the Training Data. Some subset of the training data for the adversary’s target class – i.e. ads labeled as “benign” – must include the trigger pattern.
@@
    Modify inference inputs so that they include the backdoor trigger. This is how the adversary exploits the backdoor. They overlay their chosen trigger pattern on non-benign ads, allowing those ads to circumvent the content filter and reach millions of internet users.
Note that solely achieving either of these requirements is insufficient. There will be no backdoor for the adversary to exploit if they cannot modify the training data. Similarly, the attack will fail if the adversary successfully modifies the training data but cannot submit ads which include the trigger. For example, because the adversary is not allowed, or cannot afford, to purchase the ad space. Table I and the top card on the bottom left of Figure 6 depict the requirements to perform a backdoor attack. The rest of this section describes how those requirements were determined.
The Adversarial AI literature has identified many techniques which will allow an adversary to modify training data. Four examples are listed below. A description of the required capability has been included alongside each. The first three describe techniques which allow an adversary to change benign training data. That is, the adversary has obtained some form of Write capability and, implicitly, a Read capability. The fourth technique is of particular note because the adversary has no access to benign training data. They are unable to Read or Write over benign training data.
@@
    Capability: Permanently modify training data. Modifications must be imperceptible to avoid detection.
@@
    Manipulate training data during collection or in storage. An adversary with a foothold on a victim’s data collection infrastructure can apply the trigger to benign examples as they are collected, transformed, or encoded. Similarly, an adversary could compromise the data at rest, in storage. These modifications are permanent and persist across all future uses of the data – unless the victim is able to rolls the dataset back to a previous version.
    Capability: Permanently modify training data. Modifications must be imperceptible to avoid detection.
@@
    Manipulate training data during use. The adversary transiently applies the trigger to benign data as it is used for training. For example, by compromising a data augmentation library [6].
    These modifications may not need to be imperceptible because augmented training data is typically not saved. Saving augmentation data would essentially require saving a modified version of the dataset for every single epoch of training. Doing so would incur massive storage costs while generating too much data for analysis to be feasible.
    Capability: Transiently modify training data.
@@
    Capability: Contribute training data. Modifications must be imperceptible to avoid detection.
Table I summarizes the relevant assets and minimum capabilities an adversary must obtain over each. Critically, obtaining stronger capabilities enables an adversary to perform more powerful variants of backdoor attacks. Using the minimum capabilities during their analysis allows a defender to identify the full range of backdoor attacks which are in-scope for their system. Therefore, the minimum requirements are used when mapping threats to $AF$ in Figure 6.
TABLE I: This table summarizes Backdoor Attacks. Row 3 identifies the asset-centric attack requirements for performing a backdoor attack.
@@
Attack Requirements Adversary must compromise the Training Dataset and Inference Inputs. 1. Training Dataset: Add to the training data. 2. Inference Inputs: Modify inputs so that they include the backdoor trigger.
Impact AI Output becomes adversary controlled, but only for inputs which include the backdoor trigger.
V-D Mapping an attack into a product (Phase 2)
The second step in phase 2 (4, line 4) compares the aggregated capabilities in $AF$ against the threats in $KB$ to identify in-scope threats. Having broken down an AI attack as described in Section V-C and added it into the knowledge base $KB$, defenders determine whether that attack could be enabled by obtaining any combination(s) of the capabilities identified in $AF$. As depicted in 24, line 24, this is done by comparing the capabilities an adversary could gain within the AI’s lifecycle against the attack’s requirements. If the attack’s requirements can be met by some combination of vulnerabilities, the attack is deemed to be in scope and added into $TM$. This analysis is caricatured by the red arrows in Figure 6, The defender may identify multiple vulnerabilities, across the stack and the distributed pipeline, which provide the required capability. Knowing that the adversary could exploit any of these vulnerabilities or assumptions, as a link in their chain of attacks, facilitates the defender’s risk analysis.
@@ V-D Mapping an attack into a product (Phase 2)
Notably, because $TM$ is generated using $AF$, the defender does not need in depth knowledge of each vulnerability or its originating system architecture when implementing Algorithm 3. The defender may even choose not mitigate a threat at the “root cause” vulnerabilities in $AF$. For example, a defender may opt to mitigate the risks of a backdoor attack (Table I) by implementing a consensus system instead of performing high-cost and high-complexity model retraining. This mitigation instead opts to increases confidence in AI outputs by augmenting them with information from a non-AI source – such as a LIDAR sensor in an autonomous vehicle – or a second model.
VI Enterprise RAG: A Secure Retrieval Augmented Generation Architecture

### v2

@@ III-B Training
The Training stage encompasses the processes used to create or update a Model. This stage presents unique security challenges due to the iterative nature of the training process and the complex interactions between data, model, and hyperparameters. Common operations are outlined in Figure 5. Assets present within this stage are:
  1. 1.
    Datasets: Training and Validation.
  2. 2.
    Inputs: Batched and augmented training samples.
  3. 3.
    Outputs & Output Details.
  4. 4.
    Hyperparameters: Learning rates, architectures, augmentation strategies.
  5. 5.
    Model Parameters: Weights, gradients, optimizer states.
  6. 6.
    Model Validation Criteria: Performance thresholds and early stopping rules
  7. 7.
    Model Validation Results: Accuracy metrics, loss values, confusion matrices
Critical Security Insight: The training process typically requires only read access to datasets. This principle of least privilege can be enforced to prevent dataset tampering during training.
Capability Constraints During Training:
  • Dataset Access: Read-only by design - training processes should never modify source datasets
  • Model Parameters: Full write access required - the core purpose of training
  • Transient Augmentations: Temporary modifications that don’t persist to storage
As within Dataset Collection & Assembly, the set of obtainable capabilities can be inherently constrained based on which process or operation is compromised. Augmentation, for example, modifies samples from the dataset before they are input to the model by the Training Algorithm. These modifications are transient – they are not written back to the dataset. Defenders can leverage this property to implement training processes which can access and operate on the dataset but can never modify its source; follow principles of least privilege. Similar principles apply to the validation dataset.
III-C Deploy & Inference
@@
This stage describes the process for deploying a model, serving inference requests, monitoring model performance, and collecting data for future re-training. Common operations are outlined in Figure 6. Assets present within this stage are:
  1. 1.
    Dataset: Training (implicitly, through model).
  2. 2.
    Inputs.
  3. 3.
    Outputs & Output Details.
  4. 4.
    Hyperparameters: Model, Data selection.
  5. 5.
    Model Parameters: Weights, Activations.
  6. 6.
    Monitoring Criteria.
  7. 7.
    Monitoring Results.
III-D Summary of Capabilities Across AIOps Stages
Table I summarizes how adversarial capabilities vary across AIOps stages for key asset types.
TABLE I: Adversarial Capabilities by Asset Type Across AIOps Stages
Asset Type Collection Training Deployment
Dataset Full Access Read Only Not Present
Model Params Not Present Write Required Read Only
Inputs Write (Collection) Read (Batches) Write (User)
Outputs Not Present Metrics Only Full Generation
Within Deploy & Inference, the set of capabilities which can be obtained over the model and dataset are typically constrained. Processes within this stage operate on the model but do not write it back, limiting how far integrity attacks can propagate. Adversaries cannot compromise the training dataset’s integrity, and cannot directly compromise its confidentiality, because it is not deployed alongside the model. However, the training dataset is partially encoded – memorized – within the model weights during training. This allows adversaries to indirectly inspect the training dataset by studying the model’s behavior [12, 15].
Defenders must carefully asses the flow of assets within this stage. For example, developers can reduce latency by moving the Data Collection process off the critical path for the Transmission process; i.e. for the process which responds to users or provides outputs to downstream systems. This implementation decision also limits adversarial impact to integrity or availability: Modifications within the Data Collection process would not propagate to transmitted outputs. However, this decision does not affect impact to confidentiality.
IV Challenges to Overcome
@@
Challenge 2. Understanding the impact of a vulnerability requires a holistic evaluation. Distributing an AI’s implementation and dependencies distributes its assets accordingly. Consider all the processes outlined in Figures 45, and 6. Every process can exist within a separate hardware and software context, and each performs unique operations or transformations on a subset of the system’s AI Assets. This creates many, often disjoint, attack surfaces. Reasoning about all possible implications of a vulnerability – identified in isolation, within a single context – requires that defenders map out how that vulnerability can be chained with vulnerabilities across other contexts.
@@
Our methodology addresses these challenges using an asset-centric approach. Section V provides background on asset-centric threat modeling, and Section VI describes how we build on it using the fundamental types of AI assets introduced in Section II.
V Asset-Centric Threat Modeling
Asset-Centric threat modeling presents a methodology for defenders to reason about how adversaries can impact assets and what said impact means for the system as a whole. Asset-centric threat modeling formally defines an asset as:
@@ V Asset-Centric Threat Modeling
V-A Asset-Centric Reasoning about Threats
Although some assets are inherently valuable, others are valuable because they protect, or serve as a stepping stone towards, other assets. Adversaries typically compromise the latter, through individual exploits, as they move towards their true goal: compromising one or more assets which are inherently valuable. It is therefore useful to identify the complete set of assets – of both types – an adversary must compromise along the way to their goal. Including the minimum capabilities they must obtain over each asset.
@@ V-A Asset-Centric Reasoning about Threats
  • Cryptographic Key (Asset): Inspect (Capability, compromising confidentiality).
  • Encrypted PII Data (Asset): Inspect (Capability, compromising confidentiality).
In practice and in threat models, the vulnerabilities which provide each capability can be reasoned about – and exploited – as individual steps in a chain of attacks. Adversaries might, for example, leverage side channels in a cryptographic algorithm’s implementation to reconstruct the key. Reconstructing the key is the process through which they transform the ability to indirectly inspect it into the ability to inspect it. Section VI will describe how we leverage this property to develop our methodology.
VI Asset-Centric Threat Modeling for AI
@@
In phase one (3, line 3) defenders map $AF$ by identifying capabilities adversaries could obtain over the AI asset types relevant to their $system$. Asset types and the set of possible capabilities are introduced in Section II. Defenders reason about these capabilities based on whether they originate within product scope (Section VI-A) or are inherited from consumed assets (Section VI-B). Our analysis implements this reasoning as two steps, however, there is no strict dependence between them. They can be performed in parallel with each other and alongside existing security analysis. The set of possible capabilities is also the same for both steps, the difference between them is whether each capability originates from a vulnerability or a security assumption.
Phase two performs threat analysis (2, line 2) to generate a knowledge base of attacks $KB$, and then identifies which threats from $KB$ are relevant based on $AF$ (4, line 4). Section VI-C describes how threats are broken down into their asset-centric requirements and added into $KB$. Section VI-D describes how those threats are contextualized within a product, using $AF$, to determine if, and how, they are feasible. We have developed a threat analysis engine that automates $MapThreats(AF,KB)$ (4, line 4).
@@
Phase 1, where defenders perform security analysis to generate $AF$ for their $system$, is outlined in Algorithm 2 and described in Sections VI-A and VI-B.
VI-A Identifying Adversarial Influence within Product Scope (Phase 1)
This step guides defenders through AI-centric analysis of vulnerabilities within product scope. During this process, defenders will reason about vulnerabilities from their existing security analysis but may also identify additional related vulnerabilities, such as those described in [23, 18]. As outlined in 4, line 4-line 8, this step guides defenders through identifying the AI asset types that exist within each in-scope stage, identifying vulnerabilities which could impact each asset, quantifying the capabilities each vulnerability could provide, and aggregating them into $AF$.
Repeating this process within each stage – i.e. each system boundary – allows $AF$ to describe capabilities which could be obtained across across the entire development and deployment pipelines. This allows defenders to address Challenge 1 (Section IV). An example output from this step is depicted on the top right of Figure 7.
VI-A1 Holistic Evaluation
@@ VI-A2 Enabling Collaboration
VI-A3 Reasoning about Different Adversaries
@@
VI-B Quantifying Assumptions about Adversarial Influence Outside Product Scope (Phase 1)
@@
VI-C Threat Analysis (Phase 2)
The first step in phase two (2, line 2) generates $KB$ – a prerequisite for mapping attacks (4, line 4). Defenders repeat the first portion of the procedure in 15, line 15 for every attack they wish to add into $KB$. Fortunately, this process is product-independent. $KB$ can be reused when analyzing other products because is not dependent on $system$. The cards on the bottom left of Figure 7 represent a knowledge base of threats.
As described in Section V-A and depicted in 16, line 16, an attack’s requirements can be described using the set of assets an adversary must gain capabilities over. Papernot et. al.’s work in [25] presented a first step towards describing adversarial AI attacks based on their requirements but Section II formalizes and standardizes asset types and capability ranges to the extent required for bottom-up analysis. Section VI-C1 describes the process for mapping the requirements for Backdoor attacks. Table II summarizes the identified requirements.
VI-C1 Example Breakdown: Backdoor attacks
@@
    Manipulate the Training Data, compromising its integrity. Some subset of the training data for the adversary’s target class – i.e. ads labeled as “benign” – must include the trigger pattern.
@@
    Modify inference inputs so that they include the backdoor trigger, compromising their integrity. This is how the adversary exploits the backdoor. They overlay their chosen trigger pattern on non-benign ads, allowing those ads to circumvent the content filter and reach millions of internet users.
Note that solely achieving either of these requirements is insufficient. There will be no backdoor for the adversary to exploit if they cannot modify the training data. Similarly, the attack will fail if the adversary successfully modifies the training data but cannot submit ads which include the trigger. For example, because the adversary is not allowed, or cannot afford, to purchase the ad space. Table II and the top card on the bottom left of Figure 7 depict the requirements to perform a backdoor attack. The rest of this section describes how those requirements were determined for the training data asset type.
The Adversarial AI literature has identified many techniques which will allow an adversary to compromise the training data’s integrity. Four examples are listed below. A description of the required capability has been included alongside each. The first three describe techniques which allow an adversary to change benign training data. That is, the adversary has obtained an Integrity capability and, implicitly, a Confidentiality capability. The fourth technique is of particular note because the adversary has no access to benign training data: They lack all Integrity or Confidentiality capabilities over benign training data.
@@
    Capability: Make Limited Changes. The adversary permanently modifies training data but modifications must be imperceptible to avoid detection.
@@
    Manipulate training data during collection or in storage. An adversary with a foothold on a victim’s data collection infrastructure can apply the trigger to benign examples as they are collected, transformed, or encoded. Similarly, an adversary could compromise the data at rest, in storage. These modifications are permanent and persist across all future uses of the data – unless the victim is able to roll the dataset back to a previous version.
    Capability: Make Limited Changes. The adversary permanently modifies training data but modifications must be imperceptible to avoid detection.
@@
    Manipulate training data during use. The adversary transiently applies the trigger to benign data immediately before it is used for training. For example, by compromising a data augmentation library [6].
    As augmented data is typically not collected or reviewed, adversarial changes need not be stealthy. Saving augmentation data effectively requires creating a unique copy of the dataset during every single epoch of training. Doing so incurs massive storage costs and generates too much data for analysis to be feasible.
    Capability: Make Arbitrary Changes. The adversary transiently modifies training data, which will not be reviewed, before it’s consumed by the training algorithm.
@@
    Capability: Contribute. The adversary adds to the training data but modifications must be imperceptible to avoid detection; e.g. during labeling.
Table II summarizes the relevant assets and minimum capabilities an adversary must obtain over each. Critically, obtaining stronger capabilities enables an adversary to perform more powerful variants of backdoor attacks. Using the minimum capabilities during their analysis allows a defender to identify the full range of backdoor attacks which are in-scope for their system. Therefore, the minimum requirements are used when mapping threats to $AF$ in Figure 7.
TABLE II: This table summarizes Backdoor Attacks. Row 3 identifies the asset-centric attack requirements for performing a backdoor attack.
@@
Attack Requirements Adversary must compromise the Training Dataset and Inference Inputs. 1. Training Dataset: Contribute. Adding malicious data to the training set. 2. Inference Inputs: Make Limited Changes. Adding the backdoor trigger to inputs, leaving them otherwise unchanged.
Impact Adversary can Influence Output Classification but only for inputs which include the backdoor trigger and only to convert it into the target classification.
VI-D Mapping an attack into a product (Phase 2)
The second step in phase 2 (4, line 4) compares the aggregated capabilities in $AF$ against the threats in $KB$ to identify in-scope threats. Having broken down an AI attack as described in Section VI-C and added it into the knowledge base $KB$, defenders determine whether that attack could be enabled by obtaining any combination(s) of the capabilities identified in $AF$. As depicted in 24, line 24, this is done by comparing the capabilities an adversary could gain within the AI’s lifecycle – including implied capabilities (Section II-E1) – against the attack requirements. If the attack’s requirements can be met by some combination of vulnerabilities, the attack is deemed to be in scope and added into $TM$. This analysis is caricatured by the red arrows in Figure 7, The defender may identify multiple vulnerabilities, across the stack and the distributed pipeline, which provide the required capability. Knowing that the adversary could exploit any of these vulnerabilities or assumptions, as a link in their chain of attacks, facilitates the defender’s risk analysis.
@@ VI-D Mapping an attack into a product (Phase 2)
Notably, because $TM$ is generated using $AF$, the defender does not need in depth knowledge of each vulnerability or its originating system architecture when implementing Algorithm 3. The defender may even choose not mitigate a threat at the “root cause” vulnerabilities in $AF$. For example, a defender may opt to mitigate the risks of a backdoor attack (Table II) by implementing a consensus system instead of performing high-cost and high-complexity model retraining. This mitigation instead opts to increases confidence in AI outputs by augmenting them with information from a non-AI source – such as a LIDAR sensor in an autonomous vehicle – or a second model.
VII Enterprise RAG: A Secure Retrieval Augmented Generation Architecture